### v1

Abstract
Neural networks with piecewise linear activation functions, such as rectified linear units (ReLU) or maxout, are among the most fundamental models in modern machine learning. We make a step towards proving lower bounds on the size of such neural networks by linking their representative capabilities to the notion of the extension complexity $\operatorname{xc}(P)$ of a polytope $P$, a well-studied quantity in combinatorial optimization and polyhedral geometry. To this end, we propose the notion of virtual extension complexity $\operatorname{vxc}(P)=\min\{\operatorname{xc}(Q)+\operatorname{xc}(R)\mid P+Q=R\}$. This generalizes $\operatorname{xc}(P)$ and describes the number of inequalities needed to represent the linear optimization problem over $P$ as a difference of two linear programs. We prove that $\operatorname{vxc}(P)$ is a lower bound on the size of a neural network that optimizes over $P$. While it remains open to derive strong lower bounds on virtual extension complexity, we show that powerful results on the ordinary extension complexity can be converted into lower bounds for monotone neural networks, that is, neural networks with only nonnegative weights. Furthermore, we show that one can efficiently optimize over a polytope $P$ using a small virtual extended formulation. We therefore believe that virtual extension complexity deserves to be studied independently from neural networks, just like the ordinary extension complexity. As a first step in this direction, we derive an example showing that extension complexity can go down under Minkowski sum.
Keywords:
Neural Networks Polyhedral Combinatorics Linear Optimization Minkowski Sum Extension Complexity.
1 Introduction
While machine learning is nowadays driven by rapid empirical progress, our work focuses on the theoretical understanding of the underlying models and methods. A crucial building block are feedforward neural networks. These are directed, acyclic graphs in which each vertex (neuron) defines a simple computation, usually a linear transformation composed with a scalar-valued, continuous and piecewise linear (CPWL) activation function. While a standard choice for the activation function is the rectified linear unit (ReLU) $x\mapsto\max\{0,x\}$, in this paper we focus on the more general maxout networks. These allow to compute the maximum of constantly many linear functions at each neuron. As a result, the entire network computes a (potentially complex) CPWL function. One of the big challenges in the theoretical analysis of neural networks is to understand how many neurons one requires to exactly or approximately represent a given (CPWL) function. To the best of our knowledge, it is an open question whether there exists a family of CPWL functions, which we can evaluate in polynomial time, but which cannot be represented by polynomial-size neural networks.
The piecewise linear nature of the studied networks suggests to tackle such questions by means of polyhedral geometry, see, e.g., the recent survey [25]. In fact, a similar problem to the question above used to be open for a long time in the context of linear programming, until Rothvoß [37] resolved it affirmatively: does there exist a polytope $P$ over which we can optimize in polynomial time, but any linear programming formulation must have exponential size? This question can be formalized with the notion of extension complexity $\operatorname{xc}(P)$, which describes the minimal number of facets of any polytope $Q$ that projects onto $P$. In this case, we call $Q$ an extended formulation of $P$. Rothvoß [37] proved that the matching polytope has exponential extension complexity even though Edmonds’ algorithm [11] can be used to find the maximum weight matching or minimum weight perfect matching in polynomial time. Additionally, it was also proven that a couple of polytopes associated with NP-hard optimization problems like the travelling salesperson problem have exponential extension complexity [14].
There is a direct translation between polytopes (as feasible sets of linear programs) and CPWL functions (represented by neural networks) through the notion of the support function $f_{P}(c)=\max_{x\in P}c^{\top}x$ of a polytope $P$. This (convex) CPWL function, which has one linear region for each vertex of $P$, uniquely determines $P$ via convex duality. Computing $f_{P}(c)$ means determining the objective value when optimizing over $P$ in $c$-direction. Each CPWL function $f_{P}$ can be represented by a neural network [1], though the required number of neurons can be large. To quantify this, we define the neural network complexity $\operatorname{nnc}(P)$ as the minimum number of neurons to represent $f_{P}$ by a maxout neural network.
1.1 Our Contributions
The aim of this paper is to connect the world of extended formulations with the study of neural networks. A “dream result” in this direction would be to bound $\operatorname{xc}(P)$ polynomially in $\operatorname{nnc}(P)$. Then the breakthrough results on extension complexity would directly imply strong lower bounds on the size of neural networks. It turns out, however, that there is one feature of neural networks that seems to make the “dream result” difficult to obtain: namely the ability to use subtraction. To remedy this, we propose the notion of virtual extension complexity
$\operatorname{vxc}(P)=\min\{\operatorname{xc}(Q)+\operatorname{xc}(R)\mid\text{$Q$ and $R$ are polytopes with }P+Q=R\}\ ,$
where $P+Q=\{p+q\mid p\in P,q\in Q\}$ is the Minkowski sum. In this definition, $P$ is a (formal) Minkowski difference of two polytopes $Q$ and $R$. Note that this is not the same as $R+(-1)\cdot Q$ but rather the inverse operation of Minkowski addition. The name virtual extension complexity is derived from virtual polytopes [33], a framework for the algebraic study of formal Minkowski differences of polytopes. Observe that $\operatorname{vxc}(P)\leq\operatorname{xc}(P)$ because we can always choose $Q=\{0\}$ with $\operatorname{xc}(\{0\})=0$. In that sense, virtual extension complexity is really a strengthening of the ordinary extension complexity.
In Section 3, we deduce that $\operatorname{vxc}(P)$ is indeed a lower bound for $\operatorname{nnc}(P)$, up to a constant factor. This leaves the open question to find ways to lower-bound $\operatorname{vxc}(P)$ in order to achieve the original goal to lower-bound $\operatorname{nnc}(P)$.
A consequence of the discussion above is that, if one removes the ability to subtract within neural networks, we do indeed obtain our “dream result”, namely lower bounds through $\operatorname{xc}(P)$. The resulting model is the one of monotone neural networks, which are neural networks with only nonnegative weights. In analogy to $\operatorname{nnc}(P)$, we define the monotone neural network complexity $\operatorname{mnnc}(P)$ as the minimum number of neurons to represent $f_{P}$ by a monotone maxout neural network. In Section 4, we show how lower bounds on $\operatorname{xc}(P)$ imply lower bounds on exact and approximate representations of monotone neural networks.
Studying monotone neural networks is justified both from a theoretical and a practical perspective. From the theoretical perspective, it is a natural approach in complexity theory to prove lower bounds first for monotone models of computation, to first circumvent some additional challenges of the general case; see, e.g., [42]. From the practical perspective, it is always a good idea to incorporate prior knowledge about a problem into a machine learning algorithm to solve that problem. If one knows that the target function should be a monotone function, then it might be a good idea to constrain each unit of the neural network to compute a monotone function—this exactly corresponds to nonnegative weights. See [29] and the references therein for recent studies of monotone neural networks in the machine learning community.
Summarizing, in this paper, we discuss four ways to represent a polytope or its support function through (virtual) extended formulations and (monotone) neural networks. Figure 1 shows what we know about how the associated complexity measures $\operatorname{xc}(P)$, $\operatorname{vxc}(P)$, $\operatorname{nnc}(P)$, and $\operatorname{mnnc}(P)$ relate to each other.
$\operatorname{xc}(P)$$\operatorname{mnnc}(P)$$\operatorname{vxc}(P)$$\operatorname{nnc}(P)$??
Figure 1: Relations between complexity measures for a polytope $P$. A directed arc means that the tail can be polynomially bounded by the head. It remains an open question whether $\operatorname{xc}(P)$ and $\operatorname{nnc}(P)$ can be related this way.
We conclude with some insights for $\operatorname{vxc}(P)$ and its relation to $\operatorname{xc}(P)$. Observe that $P+Q=R$ is equivalent to $f_{P}=f_{R}-f_{Q}$ pointwise. Therefore, intuitively, in order to optimize over $P$, one only needs to optimize over $R$ and $Q$ and subtract the results. We make this intuition formal in Section 5.1, implying that small extended formulations for $Q$ and $R$ are sufficient to optimize efficiently over $P$. Furthermore, in Section 5.2, we provide a class of examples with $P+Q=R$ demonstrating that $\operatorname{xc}(R)$ can be much smaller than $\operatorname{xc}(P)$. This gives important insights on how extension complexity behaves under Minkowski sum and implies that we really need to look at both $\operatorname{xc}(Q)$ and $\operatorname{xc}(R)$ in order to lower-bound $\operatorname{vxc}(P)$.
1.2 Further Related Work
In this paper we aim to prove lower bounds on neural networks representing combinatorial optimization problems. Complementing upper bounds can be found in [22] for minimum spanning trees and maximum flows. Furthermore, in [23], similar upper bounds for the knapsack problem are given, even though they are different in flavor because some integrality assumptions are made.
Concerning the general expressivity of (piecewise linear) neural networks, the celebrated universal approximation theorems state that a single layer of neurons is sufficient to approximate any continuous function on a bounded domain; see [10] for the original version for sigmoid activation functions and [27] for a version that encompasses ReLU. However, such shallow neural networks usually require a large number of neurons. A sequence of results demonstrates that deeper networks sometimes require exponentially fewer neurons to represent the same functions; see, e.g., [1, 12]. In terms of exact representation, it is known that a function can be represented if and only if it is CPWL [1], and it is still an open question whether constant depth is sufficient to do so [21, 20]. Interestingly, also for this question, monotone networks seem to be more amenable for proving lower bounds than their non-monotone counterparts [41]. Furthermore, the related question of how to efficiently write a non-convex CPWL function as a difference of two convex ones received quite some attention recently [5, 40].
We would like to emphasize that we view neural networks as a model of real-valued computation, as opposed to binary models of computation like Boolean circuits and Turing machines. In fact, if one restricts the inputs of a neural network to be binary, it is not too difficult to simulate AND-, OR-, and NOT-gates [32]. Thus, in such a binary model, every problem in P can be solved with polynomial-size neural networks. However, such networks would usually be very sensitive to single bits in the input and therefore cannot naturally be transformed into exact or approximate neural networks in the real-valued model. See also the discussion in [22] for more details.
The more useful connection to circuit complexity is through arithmetic [38], and in particular tropical circuits [26], which are also real-valued models of computation. Again we refer to [22] for a more detailed discussion of how these models relate to neural networks.
In fact, the extension complexity has been related before to Boolean and arithmetic circuits, see [13, 24]. This is also related to Goemans’ proof that the permutahedron has extension complexity $\mathcal{O}(n\log n)$ [16], as this goes via sorting networks, which can be seen as a very specific version of a piecewise-linear arithmetic circuit.
2 Preliminaries
Polyhedra and Polytopes.
A polyhedron $P$ is a finite intersection of halfspaces $P=\{x\in\mathbb{R}^{d}\mid Ax\leq b\}$; the representation as such an intersection is called $H$-representation. The affine hull of $P$ is the smallest affine subspace containing $P$ and the dimension $\dim(P)$ is defined as the dimension of its affine hull. A face $F$ of $P$ is the set of maximizers over $P$ with respect to a linear objective function: $F=\arg\max\{c^{\top}x\mid x\in P\}$ for some $c\in\mathbb{R}^{d}$. Faces of polyhedra are polyhedra themselves. Zero-dimensional faces are called vertices. Faces of dimension $\dim(P)-1$ are called facets; let $h(P)$ be the number of such facets of $P$. The minimal $H$-representation contains precisely one inequality for each facet and potentially a bunch of equalities to describe the affine hull of $P$.
If $P$ is bounded, it is also called a polytope. By the Minkowski-Weyl theorem, a polytope $P$ can equivalently be written as convex hull of finitely many points. The inclusion-wise minimal such representation is precisely the convex hull of the vertices $V(P)$, the $V$-representation. We set $v(P)=\lvert V(P)\rvert$.
Each of $V$- and $H$-representation of a polytope can be exponentially smaller than the other one. For example, the $d$-dimensional cube $\{x\in\mathbb{R}^{d}\mid\lVert x\rVert_{\infty}\leq 1\}$ has $2^{d}$ vertices but only $2d$ facets. On the other hand, the $d$-dimensional cross-polytope $\{x\in\mathbb{R}^{d}\mid\lVert x\rVert_{1}\leq 1\}$ has only $2d$ vertices, but $2^{d}$ facets. Standard references for polytope theory are [19, 45].
2.0.1 Extended Formulations.
Sometimes, a more compact way of representing a polyhedron $P$ can be obtained by representing it as the projection of a higher-dimensional polyhedron $Q\subseteq\mathbb{R}^{e}$ with $e\geq d$. Such a polytope $Q$ with $\pi(Q)=P$ for an affine projection $\pi$ is called an extended formulation of $P\subseteq\mathbb{R}^{d}$. The extension complexity of $P$ is defined as $\operatorname{xc}(P)=\min\{h(Q)\mid\pi(Q)=P\}$. The extension complexity is upper bounded by $\min\{v(P),h(P)\}$, but can be exponentially smaller. For example, the spanning tree polytope of a graph with $n$ vertices has extension complexity $\mathcal{O}(n^{3})$ [28], but the polytope itself has exponentially many vertices and facets in $n$. See [9] for a (not very recent) survey on extended formulations.
2.0.2 Support Functions.
For a polytope $P\subseteq\mathbb{R}^{d}$, its support function $f_{P}\colon\mathbb{R}^{d}\to\mathbb{R}$ is defined as $f_{P}(c)=\max_{x\in P}c^{\top}x$. In other words, it maps a linear objective direction to the optimal value obtained by optimizing over the polytope. Support functions are convex, continuous, piecewise linear, and positively homogeneous functions; here, a function $f\colon\mathbb{R}^{d}\to\mathbb{R}$ is positively homogeneous if $f(\lambda x)=\lambda f(x)$ for all scalars $\lambda\geq 0$. Let $\mathcal{F}^{d}$ be the set of all such functions from $\mathbb{R}^{d}\to\mathbb{R}$ and let $\mathcal{P}^{d}$ be the set of all polytopes embedded in $\mathbb{R}^{d}$. Then the map $\mathcal{P}^{d}\to\mathcal{F}^{d}$, $P\mapsto f_{P}$, is a bijection that is compatible with a certain set of operations:
  1. (i)
    $f_{P+Q}=f_{P}+f_{Q}$, where the “+” on the left-hand side is Minkowski sum;
  2. (ii)
    $f_{\lambda P}=\lambda f_{P}$ for $\lambda\geq 0$, where $\lambda P=\{\lambda p\mid p\in P\}$ is the dilation of $P$ by $\lambda$;
  3. (iii)
    $f_{\operatorname{conv}(P\cup Q)}=\max\{f_{P},f_{Q}\}$.
The inverse map that maps a support function $f_{P}$ to its unique associated polytope $P$ is well-studied in tropical geometry. In fact, borrowing the name from (tropical) polynomials, we call $P$ the Newton polytope of $f_{P}$. The connection between CPWL functions and their Newton polytopes has previously been used in the study of neural networks and their expressivity [44, 31, 20, 21, 6, 30].
2.0.3 Neural Networks.
For an introduction to neural networks, we refer to [17]. In this paper, we focus on so-called rank-$k$ maxout neural networks for some natural number $k\geq 2$. They generalize the well-known neural networks with rectified linear unit (ReLU) activations. Such a neural network is based on a directed, acyclic graph with node set $V$ and arcs $A\subseteq V\times V$, where the nodes are called neurons and act as computational units. The $d\geq 1$ neurons with in-degree zero are called input neurons, all other $s\geq 1$ neurons are maxout units. We assume that among the maxout units there exists exactly one neuron with out-degree zero, the output neuron. All neurons that are neither input nor output neurons are called hidden neurons. The size of the neural network is defined as the number $s$ of maxout units.
Each neuron $v$ in the neural network defines a function $z_{v}\colon\mathbb{R}^{d}\to\mathbb{R}$ as follows. Each input neuron $v$ is associated with one element $x_{i}$ of the input vector $x\in\mathbb{R}^{d}$ and simply outputs $z_{v}(x)=x_{i}$. Each maxout unit $v$ comes with a tuple of weights $w_{uv}^{i}\in\mathbb{R}$ for $i=1,\dots,k$ and all $u\in\delta^{\mathrm{in}}_{v}$, where $\delta^{\mathrm{in}}_{v}$ is the set of in-neighbors of $v$. They are naturally associated with $k$ weight vectors $w_{*v}^{i}=(w_{uv}^{i})_{u\in\delta^{\mathrm{in}}_{v}}$ of length $\lvert\delta^{\mathrm{in}}_{v}\rvert$ for $i=1,\dots,k$. The maxout neuron $v$ represents the following expression dependent on the outputs of its in-neighbors:
$z_{v}=\max\Big{\{}\sum_{u\in\delta^{\mathrm{in}}_{v}}w_{uv}^{i}z_{u}\Bigm{|}i=1,\dots,k\Big{\}}\ ,$
where $k$ is called the rank of the maxout unit.
Finally, the output of the network is defined to be the output $z_{v}$ for the output neuron $v$. Note that our neural network represents a scalar-valued function as we assumed that there is only a single output neuron. We call a maxout network monotone if all weights of all maxout neurons are nonnegative.
For a polytope $P$, we define the neural network complexity $\operatorname{nnc}(P)$ and the monotone neural network complexity $\operatorname{mnnc}(P)$ as the minimum size of a (monotone) rank-$2$ maxout network representing $f_{P}$.
Remark 1 (biases).
In practice, the units usually also involve biases. In the following, we will mainly focus on support functions of polytopes, which are positively homogeneous. Therefore, we can omit the bias in the definition of neural networks without loss of generality, compare [21, Proposition 2.3]. With our definition of a maxout network, a function $f\colon\mathbb{R}^{d}\to\mathbb{R}$ can be represented by such a network if and only if it is CPWL and positively homogeneous.
Remark 2 (maxout ranks and ReLU).
We sometimes focus on rank-2 maxout networks. To justify this, note that a rank-$k$ maxout unit can be simulated with $k-1$ rank-2 maxout units, which allows to transfer size bounds to other ranks if $k$ is viewed as a fixed constant. Moreover, observe that a rank-2 maxout network can be simulated by three parallel ReLU units because $\max\{x,y\}=\max\{0,x-y\}+\max\{0,y\}-\max\{0,-y\}$. Note, however, that this introduces negative weights to the ReLU network even if the maxout network was monotone. Therefore, unlike in the non-monotone case, monotone ReLU networks are indeed weaker than monotone maxout networks.
3 Neural Networks are Virtual Extended Formulations
In this section we prove the following theorem stating that neural network sizes can be lower-bounded through virtual extension complexity. Here and in the following, we defer some of the proofs to Appendix 0.A.
Theorem 3.1.
Let $P$ be a polytope. Then $\operatorname{vxc}(P)\leq 4\cdot\operatorname{nnc}(P)$.
We prove this theorem in three steps, represented by the next three propositions. The first proposition shows that every maxout network can be written as the difference of two monotone maxout networks each of which has the same size as the original network. This is related to writing functions represented by neural networks as tropical rational functions [44, 6].
Proposition 1.
If $f\colon\mathbb{R}^{d}\to\mathbb{R}$ is represented by a rank-$k$ maxout network of size $s$, then it can be written as a difference $f=g-h$ of two functions $g$ and $h$ that are both representable with a monotone rank-$k$ maxout network of size $s$.
The second proposition shows that small monotone neural networks imply small extended formulations of the epigraph. For a convex CPWL function, the epigraph is the polyhedron $\operatorname{epi}(f)=\{(x,t)\in\mathbb{R}^{n}\times\mathbb{R}\mid t\geq f(x)\}$. The proposition is related to earlier works connecting the extension complexity to Boolean and arithmetic circuit complexity, see e.g. [13, 24].
Proposition 2.
If $f\colon\mathbb{R}^{d}\to\mathbb{R}$ is represented by a monotone rank-$k$ maxout network of size $s\geq 1$, then $\operatorname{xc}(\operatorname{epi}(f))\leq ks$.
Proof.
Based on the neural network, we construct an extended formulation for $\operatorname{epi}(f)$ of size $ks$. For each neuron $v$, we introduce a variable $y_{v}$. For ease of notation, we identify $y_{v}$ for an input neuron $v$ with the corresponding entry of the input vector $x$. Similarly, if $v$ is the output neuron, we identify $y_{v}$ with $t$. Then, the extended formulation for the epigraph is given by the system
$y_{v}\geq\sum_{u\in\delta^{\mathrm{in}}_{v}}w_{uv}^{i}y_{u}\quad\forall\ i=1,\dots,k\quad\forall\text{ maxout units }v\enspace.$
We claim that this formulation consisting of $ks$ inequalities exactly describes $\operatorname{epi}(f)$. The crucial ingredient is the equivalence
$y_{v}\geq\sum_{u\in\delta^{\mathrm{in}}_{v}}w_{uv}^{i}y_{u}\quad\forall\ i=1,\dots,k\quad\Leftrightarrow\quad y_{v}\geq\max_{i=1,\dots,k}\sum_{u\in\delta^{\mathrm{in}}_{v}}w_{uv}^{i}y_{u}\enspace.$
Indeed, if $(x,t)\in\operatorname{epi}(f)$, setting $y_{v}\coloneqq z_{v}(x)$ to the value represented by each hidden neuron $v$ yields by construction a feasible solution with respect to all inequalities. Conversely, given a feasible solution $(x,t,y)$, one can inductively propagate the validity of the inequality $y_{v}\geq z_{v}(x)$ along a topological order of the neurons. In particular, for the output neuron, we obtain $t\geq f(x)$ and thus $(x,t)\in\operatorname{epi}(f)$. ∎
Remark 3.
In the literature, sometimes maxout networks with different maxout ranks at different units are considered. Suppose each maxout unit $v$ has its own rank $k_{v}\geq 2$. It is not difficult to see that the proof above generalizes to obtain a bound of $\sum_{v}k_{v}$ on the extension complexity of the epigraph.
The final ingredient for proving Theorem 3.1 is that the extension complexity of a polytope equals the extension complexity of the epigraph of its support function. This is related to other previous results about the extension complexity being (almost) preserved under various notions of duality, see, e.g., [28, 18, 43].
Proposition 3.
It holds that $\operatorname{xc}(P)=\operatorname{xc}(\operatorname{epi}(f_{P}))$ for all polytopes $P$ with $\dim(P)\geq 1$.
Proof of Theorem 3.1.
The statement simply follows from combining Propositions 1, 2 and 3. The factor $4$ arises from one factor $2$ through Proposition 1 and another factor $2$ from Proposition 2 as $\operatorname{nnc}(P)$ is defined with $k=2$. ∎
4 Implications for Monotone Neural Networks
The ultimate goal of connecting neural networks and extended formulations in this paper is to use existing powerful results for lower bounds on the extension complexity to obtain lower bounds on neural networks. To pursue this for general neural networks though, we would need lower bounds on the (potentially stronger) notion of virtual extension complexity, which are currently unavailable. However, if we restrict our attention to monotone neural networks, lower bounds on the extension complexity immediately translate to lower bounds on this subclass of neural networks, as we demonstrate.
Theorem 4.1.
Let $P$ be a polytope. Then $\operatorname{xc}(P)\leq 2\cdot\operatorname{mnnc}(P)$.
Proof.
Combine Propositions 2 and 3. The factor $2$ arises from Proposition 2 as $\operatorname{mnnc}(P)$ is defined with $k=2$. ∎
In particular, we obtain the following bounds on monotone neural networks based on known bounds on the extension complexity. For a complete graph $(V,E)$ with $n$ vertices and a weight vector $c=(c_{e})_{e\in E}\in\mathbb{R}^{E}$, let $f_{\mathrm{MAT}}(c)$ be the value of a maximum weight matching and $f_{\mathrm{TSP}}(c)$ the value of a longest [^1] This is to make the function convex; this is equivalent to finding the shortest tour with respect to the negated weights. traveling salesperson (TSP) tour with respect to $c$. We denote the matching polytope by $P_{\mathrm{MAT}}$ and the TSP polytope by $P_{\mathrm{TSP}}$.
Theorem 4.2.
If a monotone maxout network represents $f_{\mathrm{MAT}}$ or $f_{\mathrm{TSP}}$, then it must have size at least $2^{\Omega(n)}$.
Proof.
The functions $f_{\mathrm{MAT}}$ and $f_{\mathrm{TSP}}$ are the support functions of $P_{\mathrm{MAT}}$ and $P_{\mathrm{TSP}}$, respectively. Both have extension complexity $2^{\Omega(n)}$ [37]. Now, the result follows from Theorem 4.1. ∎
In the same way, one can prove lower bounds for neural networks computing the support function of any polytope with high extension complexity, e.g., for neural networks solving the MAX-CUT problem or the stable set problem. The corresponding lower bounds on the extension complexity were first proven by Fiorini, Massar, Pokutta, Tiwary, and de Wolf [14]. Furthermore, they also derived that the extension complexity of the TSP polytope is at least $2^{\Omega(\sqrt{n})}$, before Rothvoß [37] improved it to $2^{\Omega(n)}$.
We would like to explicitly highlight one additional result building on the following, which we will use again later.
Theorem 4.3 ([36]).
There is a family of matroids $M_{n}$ on $n$ elements whose matroid base polytopes have extension complexity $2^{\Omega(n)}$.
In fact, the proof of this theorem in [36] shows that almost all matroid base polytopes must have exponential extension complexity. With Theorem 4.1, we get the following implication. For a family of matroids $M_{n}$ on $n$ elements and a weight vector $c\in\mathbb{R}^{n}$, let $f_{n}(c)$ be the value of the maximum weight basis in $M_{n}$.
Theorem 4.4.
There exists a family of matroids $M_{n}$ on $n$ elements such that every monotone maxout neural network representing $f_{n}$ must have size $2^{\Omega(n)}$.
For machine learning applications, it is arguably less important to obtain neural networks that exactly represent a given function. Instead, approximating the desired output is often sufficient. We demonstrate that also in the approximate setting, lower bounds on the extension complexity can be transferred to neural networks. We would like to emphasize that there exist many different ways to define what it means to approximate a given function or problem “sufficiently well”, both in optimization and machine learning. The “correct” notion always depends on the context. See also [30] for a discussion of how approximations can be translated between polytopes and neural networks.
Theorem 4.5.
Suppose a monotone maxout neural network represents a function $f$ such that $f_{\mathrm{MAT}}(c)\leq f(c)\leq(1+\epsilon)f_{\mathrm{MAT}}(c)$ for all $c\in\mathbb{R}_{\geq 0}^{E}$. Then the neural network must have size at least $2^{\Omega(\min\{n,1/\epsilon\})}$.
In particular, monotone neural networks cannot serve as fully-polynomial approximation schemes for the matching problem, in the sense that they cannot have polynomial size in both $n$ and $1/\epsilon$ to approximate the matching problem to $\epsilon$-precision. We understand Theorem 4.5 as a prototype for the fact that, in principle, inapproximability results from extended formulations can be transferred to monotone neural networks. See, e.g., a framework for such lower bounds in [7] and inapproximability up to a factor $2$ for vertex cover in [2].
5 Properties of Virtual Extension Complexity
In this section, we give some insights towards understanding the concept of virtual extension complexity. First, we argue that, even though virtual extended formulations are more general than the ordinary extended formulations, they can still be used to optimize efficiently via linear programming. In the second part of the section, we give an example demonstrating that Minkowski sum can indeed drastically reduce the extension complexity. More precisely, there exist polytopes $P+Q=R$ with $\operatorname{xc}(R)$ being way smaller than $\operatorname{xc}(P)$. This means that for obtaining useful lower bounds on $\operatorname{vxc}(P)$, we indeed need to look at both, $\operatorname{xc}(Q)$ and $\operatorname{xc}(R)$, and cannot just focus on $\operatorname{xc}(R)$.
5.1 Optimizing over Virtual Extended Formulations
One of the main reasons to study extension complexity is that it quantifies how well a given problem can be formulated as a linear program. Once we have a small-size extended formulation for a polytope $P$, we can efficiently optimize over $P$ by solving a single linear program. In this section, we argue that virtual extension complexity allows a natural progression of this idea: it describes the power of differences of two linear programs. Once we have a small-size virtual extended formulation for a polytope $P$, we can optimize efficiently over $P$ by solving two linear programs.
To this end, assume that we have polytopes $P+Q=R$ with small extended formulations for $Q$ and $R$. Switching to support functions implies that for all objective directions $c\in\mathbb{R}^{d}$ we obtain $\max_{x\in P}c^{\top}x+\max_{x\in Q}c^{\top}x=\max_{x\in R}c^{\top}x$. Thus, to obtain the optimal objective value, it is sufficient to have small extended formulations for $Q$ and $R$, as we can really just optimize over $Q$ and $R$ and subtract the results. We now argue that this is basically also true for the solution itself.
Proposition 4.
Let $P+Q=R$ be polytopes and let $c\in\mathbb{R}^{d}$ be an objective direction such that the linear program $\max_{x\in R}c^{\top}x$ has a unique optimal solution $x^{R}$. Then also the linear programs $\max_{x\in P}c^{\top}x$ and $\max_{x\in Q}c^{\top}x$ have unique solutions $x^{P}$ and $x^{Q}$, respectively, and it holds that $x^{P}=x^{R}-x^{Q}$.
The proof is based on the following fact about faces of Minkowski sums, which can be found, e.g., in [19].
Lemma 1.
Let $P+Q=R$ be polytopes and $c$ be an objective direction. Then, the optimal face in $c$-direction of $R$ is the Minkowski sum of the optimal faces in $c$-direction of $P$ and $Q$, respectively.
Proposition 4 implies that, for generic objective directions $c$, we can efficiently find the optimal solution within $P$ by linear programming using small extended formulations for $Q$ and $R$. However, the situation becomes more tricky if multiple optimal solutions exist within $R$, and then potentially also within $Q$. Indeed, in this case we would not know which of the optimal solutions we should subtract in order to obtain a solution that is feasible for $P$. Therefore, we need to be more careful to handle non-generic objective functions $c$. One solution would of course be to randomly perturb the objective function $c$ by a tiny amount such that the optimal solution in $R$ becomes unique. In practice, however, it is not clear how to ensure that the perturbation is actually small enough and generic enough at the same time.
The key idea is that we need to make sure that ties between equally good solutions are broken consistently across $Q$ and $R$. Therefore, we suggest a deterministic procedure based on a lexicographic objective function. By this, we mean a sequence of objective functions where later entries are only considered to break ties.
Proposition 5.
Let $P+Q=R$ be polytopes and $(c_{1}^{\top}x,\dots,c_{d}^{\top}x)$ be a lexicographic objective function. If $C\coloneqq\{c_{1},\dots,c_{d}\}$ forms a basis of $\mathbb{R}^{d}$, the sets of maximizers for $P$, $Q$ and $R$ are singletons $\{x^{P}\},\{x^{Q}\}$ and $\{x^{R}\}$, respectively, and $x^{P}+x^{Q}=x^{R}$.
This lays the basis for Algorithm 1.
1 let $c_{1}\coloneqq c\in\mathbb{R}^{d}$;
2 compute an arbitrary basis $C\coloneqq\{c_{1},c_{2},\dots,c_{d}\}$ of $\mathbb{R}^{d}$;
3 let $x^{Q}$ and $x^{R}$ be optimal solutions within $Q$ and $R$, respectively, with respect to the lexicographic objective function $(c_{1}^{\top}x,c_{2}^{\top}x,\dots,c_{d}^{\top}x)$;
4 return $x^{R}-x^{Q}$;
Algorithm 1 Solve $\max_{x\in P}c^{\top}x$ where $P$ is given through extended formulations of two polytopes $Q$ and $R$ with $P+Q=R$.
Theorem 5.1.
Algorithm 1 correctly returns an optimal solution of $\max_{x\in P}c^{\top}x$ and can be implemented with a running time that is polynomial in the encoding sizes of $c$ and the extended formulations for $Q$ and $R$.
5.2 Extension Complexity and Minkowski Sums
For a better understanding of virtual extension complexity, the behavior of extension complexity under Minkowski sum is crucial. It is well-known that $xc(R)\leq xc(P)+xc(Q)$ for $P+Q=R$. This follows from plugging the extended formulations for $P$ and $Q$ into the definition of the Minkowski sum. However, for bounding virtual extension complexity, one would need to bound $\operatorname{xc}(P)$ in terms of $\operatorname{xc}(Q)$ and $\operatorname{xc}(R)$. In the following, we present a class of examples demonstrating that one cannot simply focus on $\operatorname{xc}(R)$ for this, but also needs to take $\operatorname{xc}(Q)$ into account.
Recall that the regular permutahedron for $n\in\mathbb{N}$ is the polytope
$\Pi_{n}=\operatorname{conv}\{(1,2,\ldots,n)^{\sigma}\mid\sigma\in S_{n}\}$
arising as the convex hull of all permutations of the vector $(1,2,\ldots,n)$. The regular permutahedron has the property that all edge directions are of the form $e_{i}-e_{j}$ for canonical unit vectors $e_{i},e_{j}$. More generally, polytopes fulfilling this form the class of generalized permutahedra. A different characterization is in terms of Minkowski sums: a polytope $P$ is a generalized permutahedron if and only if there is a polytope $Q$ with $P+Q=\lambda\cdot\Pi_{n}$ for some $\lambda\geq 0$, see [34, Proposition 3.2]. It turns out that every matroid base polytope is such a generalized permutahedron, and we can even choose $\lambda=1$ in this case [4, 15].
Hence, we can apply Theorem 4.3 to deduce a statement about summands of $\Pi_{n}$.
Theorem 5.2.
For $n\in\mathbb{N}$, there is a polytope $\Pi_{n}$ with extension complexity $\mathcal{O}(\text{poly}(n))$, such that a Minkowski summand has extension complexity $2^{\Omega(n)}$.
Proof.
By [16], the extension complexity of $\Pi_{n}$ is $\Theta(n\log(n))$. As discussed, each matroid base polytope $P$ is a Minkowski summand of $\Pi_{n}$. Now, Theorem 4.3 guarantees that there exist such polytopes with extension complexity $2^{\Omega(n)}$. ∎
Considering a matroid polytope $P$ with high extension complexity from the latter proof such that $P+Q=\Pi_{n}$ for some other generalized permutahedron $Q$, we cannot exclude that $Q$ has also high extension complexity. Therefore, without a better understanding of $Q$, this does not yield any useful bound on $\operatorname{vxc}(P)$.
On a final note, we would like to emphasize that in the definition of the virtual extension complexity, it is important that $Q$ and $R$ are polytopes and not potentially unbounded polyhedra. A simple example is given by the choice of $Q=R=\mathbb{R}^{d}$, which would immediately imply $\operatorname{vxc}(P)=0$ for any polyhedron $P$, if it was allowed. The important algebraic property that we need to make sure that we can recover $P$ from $Q$ and $R$ is the cancellation property: For polytopes $A,B,C$, it is true that $A+C=B+C$ implies $A=B$. The same is not true for potentially unbounded polyhedra.
6 Open Questions
Clearly, the most intriguing open question is: Can $\operatorname{vxc}(P)$ be (much) smaller than $\operatorname{xc}(P)$? If the answer is no, then this would imply strong lower bounds on neural networks. On the other hand, if the answer is yes, it implies that solving two linear programs and taking the difference is (much) more powerful than just solving one linear program. In fact, we are not even aware of an example for which we can prove that $\operatorname{vxc}(P)<\operatorname{xc}(P)$. On the other hand, as for the usual extension complexity, we do not expect $\operatorname{vxc}(P)$ to be polynomial for any polytope $P$ associated with an NP-hard optimization problem. An obvious, but seemingly challenging approach to lower-bound virtual extension complexity is to take any previously successful method for lower-bounding extension complexity and try to adapt it such that it also works in the virtual case.
We hope that this paper inspires researchers to answer this question: Are two LPs better than one?
Acknowledgments
We thank Alex Black, Daniel Dadush, Jesús A. De Loera, Samuel Fiorini, Neil Olver, László Végh, Matthias Walter, and Stefan Weltge for insightful discussions around (virtual) extension complexity that have shaped this paper.
Christoph Hertrich is supported by the European Union’s Horizon Europe research and innovation program under the Marie Skłodowska-Curie grant agreement No 101153187—NeurExCo.
Appendix 0.A Omitted proofs
Proof of Proposition 1.
Let $\mathcal{N}_{0}$ be the maxout network of size $s$ representing $f$. Without loss of generality, we assume that the underlying graph of $\mathcal{N}_{0}$ is transitively closed. We can always achieve this by introducing arcs with weights equal to zero. This does not increase the number of neurons and therefore not the size of the neural network. Let $v_{1},v_{2},\dots,v_{s}$ be a topological order of the maxout units of $\mathcal{N}_{0}$.
We construct a sequence of equivalent neural networks $\mathcal{N}_{1},\dots,\mathcal{N}_{s-1}$, pushing the negative weights neuron by neuron towards the output. More precisely, in $\mathcal{N}_{p}$ none of the neurons $v_{1}$ to $v_{p}$ will have negative weights on their incoming arcs.
To go from $\mathcal{N}_{p-1}$ to $\mathcal{N}_{p}$ for some $p\geq 1$, we let the first $p-1$ neurons unchanged. Now consider the weights $w_{uv}^{i}$ of $v\coloneqq v_{p}$. We split them into a positive and negative part, that is, $w_{uv}^{i}=a_{uv}^{i}-b_{uv}^{i}$ with $a_{uv}^{i}=\max\{w_{uv}^{i},0\}$, $b_{uv}^{i}=\max\{-w_{uv}^{i},0\}$. Observe that
$z_{v}$ $=\max\Big{\{}\sum_{u\in\delta^{\mathrm{in}}_{v}}\big{(}a_{uv}^{i}+\sum_{j\neq i}b_{uv}^{j}\big{)}z_{u}\Bigm{|}i=1,\dots,k\Big{\}}-\sum_{u\in\delta^{\mathrm{in}}_{v}}\sum_{j=1}^{k}b_{uv}^{j}z_{u}\enspace.$ (1)
This allows to construct $\mathcal{N}_{p}$ from $\mathcal{N}_{p-1}$ by the following operation. Each weight $w_{uv}^{i}$ of $v=v_{p}$ is set to $a_{uv}^{i}+\sum_{j\neq i}b_{uv}^{j}$. To make up for this change, one has to correct the weights of neurons $v_{q}$ with $q>p$ to incorporate the term outside the maximum in (1). More precisely, if $\widetilde{v}\coloneqq v_{q}$ is an out-neighbor and $u\in\delta^{\mathrm{in}}_{v}$ is an in-neighbor of $v=v_{p}$, then we need to update the weight $w_{u\widetilde{v}}^{i}$ to $w_{u\widetilde{v}}^{i}-w_{v\widetilde{v}}^{i}\sum_{j=1}^{k}b_{uv}^{j}$. It is easy to verify that this weight update exactly recovers the missing term from (1) when computing $z_{\widetilde{v}}$. This might introduce new negative weights, but all of them are associated with neurons in the topological order after $p$. So, after doing this operation, we obtain an equivalent neural network $\mathcal{N}_{p}$ with only nonnegative weights associated with the neurons $v_{1}$ to $v_{p}$.
We continue that way until we obtain $\mathcal{N}_{s-1}$, where all weights except those of the output neuron are nonnegative. Performing the same splitting of the weights into positive and negative parts as before and looking at the expression of (1) for the output neuron $v\coloneqq v_{s}$, we define $z_{v}\eqqcolon g-h$ according to (1). While $g$ is already a maxout expression with only nonnegative weights, observe that $h$ can be artificially converted into such a maxout expression by taking the maximum over the same linear expression in the $z_{u}$-terms $k$ times.
Hence we can obtain two monotone neural networks computing $g$ and $h$ by simply copying all neurons except the output neuron from $\mathcal{N}_{s-1}$ and using the maxout expressions described above as weights of the respective output neurons in the two neural networks. ∎
Proof of Proposition 3.
The key idea behind this statement is the well-known fact that the dual of the epigraph of $f_{P}$ as a cone in $\mathbb{R}^{d+1}$ equals the homogenization of $P$, see, e.g., [3, 35]. We give a short geometric proof of this fact to show the ingredients of the construction. A point $(c,t)$ in the epigraph $\operatorname{epi}(f_{P})$ fulfills the equivalent conditions
$t\geq f(c)\ \Leftrightarrow\ t\geq\max_{x\in P}c^{\top}x\ \Leftrightarrow\ t\geq c^{\top}x\;\forall x\in P\ \Leftrightarrow\ c^{\top}x-t\leq 0\;\forall x\in P\ .$
This yields the equivalent representation
$\operatorname{epi}(f_{P})=\Big{\{}(c,t)\in\mathbb{R}^{d}\times\mathbb{R}\Bigm{|}(c^{\top},t)\begin{pmatrix}x\\ -1\end{pmatrix}\leq 0\;\forall x\in P\Big{\}}\enspace.$
This is the polar cone of the cone over $P\times\{-1\}$ that is
$\operatorname{epi}(f_{P})=\text{cone}\left(\left\{(x,-1)\mid x\in P\right\}\right)^{\star}\enspace.$
To finish the proof, we argue that (de-)homogenization and cone polarity preserve extension complexity. The claim about (de-)homogenization follows from the discussion before [18, Theorem 6]. The statement about cone polarity is implied by the discussion before [18, Proposition 2]. ∎
Proof of Theorem 4.5.
Consider the Newton polytope $P_{f}$ of the function $f$ computed by the neural network. Note that $P_{f}$ is contained in the nonnegative orthant since the weights of a monotone neural network are nonnegative. We slightly modify $P_{f}$ to also include all nonnegative vectors that are smaller than a vector in $P_{f}$ and obtain
$Q\coloneqq\{x\in\mathbb{R}^{E}\mid\exists y\in P_{f}\colon 0\leq x\leq y\}\ .$ (2)
With that modification, $Q$ is a “monotone polytope” [^2] Note that this notion of a monotone polytope does not correspond to the notion of monotone neural networks. as defined in [37]. Moreover, we still maintain the inequality $f_{\mathrm{MAT}}(c)\leq f_{Q}(c)\leq(1+\epsilon)f_{\mathrm{MAT}}(c)$ for all $c\in\mathbb{R}_{\geq 0}^{E}$. As argued around [37, Equation 7], this is equivalent to $P_{\mathrm{MAT}}\subseteq Q\subseteq(1+\epsilon)P_{\mathrm{MAT}}$. By [37, Corollary 4.2], this implies that $\operatorname{xc}(Q)\geq 2^{\Omega(\min\{n,1/\epsilon\})}$; see also [39, 8]. The definition (2) of $Q$ implies $\operatorname{xc}(Q)\leq\operatorname{xc}(P_{f})+2\cdot\lvert E\rvert\leq\operatorname{xc}(P_{f})+\mathcal{O}(n^{2})$. Thus, we obtain the same asymptotic lower bound on $\operatorname{xc}(P_{f})$. The statement then follows by Theorem 4.1. ∎
Proof of Proposition 4.
As for $R$ the optimal face in $c$-direction is the singleton $\{x^{R}\}$, by Lemma 1 the corresponding faces of $P$ and $Q$ must be singletons, too, namely $\{x^{P}\}$ and $\{x^{Q}\}$ with $x^{P}+x^{Q}=x^{R}$. ∎
Proof of Proposition 5.
For $i=1,\dots,d$, let $F^{P}_{i}$ be the set of optimal solutions over $P$ with respect to the lexicographic objective function $(c_{1}^{\top}x,\dots,c_{i}^{\top}x)$. Let $F^{Q}_{i}$ and $F^{R}_{i}$ be defined analogously. We show by induction on $i$ that $F^{P}_{i}+F^{Q}_{i}=F^{R}_{i}$. For $i=1$, this follows from Lemma 1. For $i>1$, observe that $F^{P}_{i}$ is the face in $c_{i}$-direction of $F^{P}_{i-1}$, and analogously for $Q$ and $R$. Hence, we can apply Lemma 1 again to complete the induction step. As a consequence, we obtain that $F^{P}_{d}+F^{Q}_{d}=F^{R}_{d}$. Moreover, since $C$ is a basis of $\mathbb{R}^{d}$, it follows that any two optimal solutions with respect to the lexicographic objective function must be equal, implying that $F^{P}_{d}$, $F^{Q}_{d}$, and $F^{R}_{d}$ are singletons. ∎
Proof of Theorem 5.1.
Let $x^{P},x^{Q},x^{R}$ be the unique optimizers as guaranteed by Proposition 5. As they fulfill $x^{P}+x^{Q}=x^{R}$, the algorithm correctly returns $x^{P}=x^{R}-x^{Q}$. Furthermore, since $x^{P}$ is optimal with respect to the lexicographic objective, it is in particular optimal with respect to the first component of the objective, which is $c^{\top}x$. Therefore, the algorithm returns an optimal solution for $\max_{x\in P}c^{\top}x$.
For the running time, the only interesting step is line 1. One way to see that this runs in polynomial time is to first solve with respect to objective $c_{1}$, obtain the optimal objective value $\lambda_{1}$, then add the constraint $c_{1}^{\top}x=\lambda_{1}$, and solve with respect to objective $c_{2}$, and so on. After solving $d$ linear programs for each of the two polytopes $Q$ and $R$, we obtain the desired optimal solutions with respect to the lexicographic objective. Depending on the precise linear programming algorithm used, one can handle the lexicographic objective function directly in a single linear programming computation for each of the two polytopes. ∎